# Dimensions underlying the representational alignment of deep neural networks with humans

### v1

@@ Abstract
Determining the similarities and differences between humans and artificial intelligence is an important goal both in machine learning and cognitive neuroscience. However, similarities in representations only inform us about the degree of alignment, not the factors that determine it. Drawing upon recent developments in cognitive science, we propose a generic framework for yielding comparable representations in humans and deep neural networks (DNN). Applying this framework to humans and a DNN model of natural images revealed a low-dimensional DNN embedding of both visual and semantic dimensions. In contrast to humans, DNNs exhibited a clear dominance of visual over semantic features, indicating divergent strategies for representing images. While in-silico experiments showed seemingly-consistent interpretability of DNN dimensions, a direct comparison between human and DNN representations revealed substantial differences in how they process images. By making representations directly comparable, our results reveal important challenges for representational alignment, offering a means for improving their comparability.
Today’s deep neural networks (DNNs) have shown remarkable performance, offering powerful models that often approach or even exceed human performance across diverse perceptual and cognitive benchmarks. Paralleling their success in machine learning, recent work in computational cognitive neuroscience has demonstrated striking similarities between artificial and biological processing systems (Güçlü and van Gerven, 2015; Khaligh-Razavi and Kriegeskorte, 2014; Yamins et al., 2014; Kubilius et al., 2016; Rajalingham et al., 2015). This has sparked considerable interest in linking DNN representations and behaviors to those found in humans. From the machine learning perspective, understanding the limitations of DNNs can support the development of artificial intelligence systems that are better aligned with humans, promising improved and more robust performance (Sucholutsky et al., 2023). From the computational cognitive neuroscience perspective, DNNs with stronger human alignment promise to be better candidate computational models of human cognition and behavior (Cichy and Kaiser, 2019; Lindsay, 2021; Kanwisher et al., 2023; Doerig et al., 2023).
@@
Despite the appeal of global metrics for comparing the representational alignment of humans and DNNs, they only provide a quantification of the degree of representational or behavioral alignment. However, without explicit hypotheses about potential causes for misalignment, global metrics are limited in their explanatory scope of what features determine this degree of alignment, that is, what representational factors underlie the similarities and differences between humans and DNNs. While diverse methods for interpreting DNN activations have been developed at various levels of analysis, ranging from single units to entire layers (Bau et al., 2020; Zhou et al., 2018; Morcos et al., 2018), the direct comparability to human representations has remained a key challenge.
Inspired by recent work in the cognitive sciences that has revealed core visual and semantic representational dimensions underlying human similarity judgments of object images (Hebart et al., 2020), here we propose a framework for systematically analyzing and comparing the dimensions underlying the representations in DNNs and humans. In this work, we apply this framework to human visual similarity judgments and representations in a DNN trained to classify natural images. Our approach reveals numerous interpretable DNN dimensions that appear to reflect both visual and semantic image properties and that appear to be well-aligned to humans. In contrast to humans who showed a dominance of semantic over visual dimensions, DNNs exhibited a striking visual bias, which only emulates human semantic behavior. While psychophysical experiments on DNN dimensions underscored their global interpretability, a face-to-face comparison of dimensions with humans revealed that DNN representations in fact only approximate human representations but lack the consistency expected from feature-specific visual and semantic dimensions. Together, our results reveal key factors underlying the representational alignment and misalignment between humans and DNNs, shed light on potentially divergent representational strategies, and highlight the potential of this approach for identifying determinants of the similarities and differences between both domains.
@@ 1 Results
Figure 1: Overview: A computational framework that captures core DNN object representations in analogy to humans by simulating behavioral decisions in an odd-one-out task. a, The triplet odd-one-out task, where a human participant or a DNN is presented a set of three images and is asked to select the image that is most different from the others. b, Sampling approach of odd-one-out decisions from DNN representations. First, a dot-product similarity space is constructed from DNN features. Next, for a given triplet of objects, the most similar pair in this similarity space is identified, making the remaining object the odd-one-out. c, Illustration of the computational modeling approach to learn a lower-dimensional object representation for human participants and the DNN, optimized to predict behavioral choices made in the triplet task. d, Schematic depiction of the interpretability pipeline that allows predicting object embeddings from pretrained DNN features.
To improve the comparability of human and DNN representations, we aimed at identifying the similarities and differences in key dimensions underlying their image representations. To achieve this aim, we treated the neural network analogous to a human participant carrying out a cognitive behavioral experiment and then derived representational embeddings both from human similarity judgments and a DNN on the same behavioral task. This approach ensured direct comparability between human and DNN representations. As a behavioral task, we chose a triplet odd-one-out similarity task, where from a set of three object images $i$, $j$, $k$ a participant has to select the most dissimilar object (Fig. 1a). In this task, the perceived similarity between two images $i$ and $j$ is defined as the probability of choosing these images to belong together across varying contexts imposed by a third object image $k$. By virtue of providing minimal contexts, the odd-one-out task highlights the information sufficient for capturing the similarity between object images $i$ and $j$ across diverse contexts. In addition, it approximates human categorization behavior for arbitrary visual and semantic categories, even for quite diverse sets of objects (Zheng et al., 2019; Hebart et al., 2020; Muttenthaler et al., 2022). Thus, by focusing on the building blocks of categorization which underlies diverse behaviors, this task is ideally suited for comparing object representations between humans and DNNs.
@@
For humans, we used a set of 4.7 million publicly available odd-one-out judgments (Hebart et al., 2023) on 1,854 diverse object images, derived from the THINGS object concept and image database (Hebart et al., 2019). For the DNN, we collected similarity judgments for 24,102 images of the same objects used for humans (1,854 objects, 13 examples per object). We used a larger set of object images since the DNN was less limited by constraints in dataset size than humans. This allowed us to obtain more precise estimates of their representation. As a DNN, we chose a pretrained VGG-16 model (Simonyan and Zisserman, 2014), given its common use in the computational cognitive neurosciences. Specifically, this network has been shown to exhibit good correspondence to both human behavior (Geirhos et al., 2018) and measured neural activity (Güçlü and van Gerven, 2015; Schrimpf et al., 2018; Nonaka et al., 2021) and performs well at predicting human similarity judgments (Jozwik et al., 2017; Peterson et al., 2018; King et al., 2019; Storrs et al., 2021; Kaniuth and Hebart, 2022; Muttenthaler et al., 2023a). However, for completeness, we additionally ran similar analyses for a broader range of neural network architectures (see Supplementary Information A). We focused on the penultimate layer activations as they reflect the most high-level abstraction of the input and are thus representationally closer to the behavioral outputs. For the DNN, we generated a dataset of behavioral odd-one-out choices for the 24,102 object images (Fig. 1b). To this end, we first extracted the DNN layer activations for all images. Next, for a given triplet of activations ${\bm{z}}_{i}$, ${\bm{z}}_{j}$ and ${\bm{z}}_{k}$, we computed the dot product between each pair as a measure of similarity, then identified the most similar pair of images in this triplet, and designated the remaining third image as the odd-one-out. Given the excessively large number of possible triplets for all 24,102 images, we approximated the full set of object choices from a random subset of $20$ million triplets (Jain et al., 2016).
@@
From both sets of available triplet choices, we next generated two representational embeddings, one for humans and one for the DNN. In these embeddings, each object is characterized by a set of representational dimensions. The embeddings were optimized for predicting the odd-one-out choices in humans and DNNs, respectively. For comparability to previous work in humans (Zheng et al., 2019; Hebart et al., 2020; Muttenthaler et al., 2022), we imposed sparsity and non-negativity constraints on the optimization, which support their interpretability and provide cognitive plausible criteria for dimensions(Hebart et al., 2020, 2023; Hoyer, 2002; Murphy et al., 2012; Fyshe et al., 2015). Sparsity limited the number of dimensions used, while non-negativity ensured dimensions only added to the explanation of behavior without canceling each other out. During training, each randomly initialized embedding was optimized using a recent variational embedding technique (Muttenthaler et al., 2022) (see Methods for details). The optimization resulted in two stable, low-dimensional embeddings, with 70 reproducible dimensions for the DNN embedding and 68 for the human embedding. The DNN embedding captured 84.03% of the total variance in image-to-image similarity, while the human embedding captured 90.85% of the explainable variance given the empirical noise ceiling of the dataset.
1.1 DNN dimensions reflect diverse conceptual and perceptual properties
@@
Having identified stable, low-dimensional embeddings that are predictive of triplet odd-one-out judgments, we first assessed the interpretability of each identified DNN dimension by visualizing object images with large numeric weights. In addition to this qualitative assessment, we validated these observations for the DNN by asking 12 (6 female, 6 male) human participants to provide labels for each dimension separately. Similar to the core semantic and visual dimensions underlying odd-one-out judgments in humans described previously (Hebart et al., 2020; Muttenthaler et al., 2022; Hebart et al., 2023), the DNN embedding yielded many interpretable dimensions, which appeared to reflect both semantic and visual properties of objects. The semantic dimensions included taxonomic membership (e.g. food-related, technology-related, home-related) and other knowledge-related features (e.g. softness), while the visual dimensions reflected visual-perceptual attributes (e.g. round, green, stringy), with some dimensions reflecting a composite of semantic and visual features (e.g. green and organic) (Fig. 2a). Of note, the DNN dimensions also revealed a sensitivity to basic shapes, including roundness, boxiness and tube-shape. This suggests that, in line with earlier studies (Hermann et al., 2020; Singer et al., 2022), DNNs indeed learn to represent basic shape features, an aspect that might not be apparent in their overt behavior (Geirhos et al., 2019).
@@
Despite the apparent similarities, there were, however, also striking differences found between humans and the DNN. First, overall, DNN dimensions were less interpretable than human dimensions, as confirmed by the evaluation of all dimensions by two independent raters (see Supplementary Information B). This indicates a global difference in how the DNN assigns images as being conceptually similar to each other. Second, while human dimensions were clearly dominated by semantic properties, many DNN dimensions were more visual-perceptual in nature or reflected a mixture of visual and semantic information. We quantified this observation by asking the same two independent experts to rate human and DNN dimensions according to whether they were primarily visual-perceptual, semantic, reflected a mixture of both, or were unclear (Fig. 2b). To confirm that the results were not an arbitrary byproduct of the chosen DNN architecture, we provided the raters with four additional DNNs for which we had computed additional representational embeddings. The results revealed a clear dominance of semantic dimensions in humans, with only a small number of mixed dimensions. In contrast, for DNNs we found a consistently larger proportion of dimensions that were dominated by visual information or that reflected a mixture of both visual and semantic information (Fig. 2c, Supplemental Fig. S1b) for all DNNs). This demonstrates a clear difference in the relative weight that humans and DNNs assign to visual and semantic information, respectively.
@@ 1.2 Linking DNN dimensions to their interpretability
Figure 3: Relevance of image features for embedding dimension. a, General methodology of the approach. We used Grad-CAM (Selvaraju et al., 2017) to visualize the importance of distinct image parts based on the gradients of the penultimate DNN features that we initially used to sample triplet choices. The gradients were obtained in our fully differentiable interpretability model with respect to a dimension ${\bm{w}}$ in our embedding. b, We visualize the heatmaps for three different images and dimensions. Each column shows the relevance of parts of an image for that dimension. For this figure, we filtered the embedding by images available in the public domain. Images used under a CC0 license, from Flickr: Cezary Borysiuk, Wojtek Szkutnik.
@@
Despite the overall differences in human and DNN representational dimensions, the DNN also contained many dimensions that appeared to be interpretable and comparable to those found in humans. Next, we aimed at testing to what degree these interpretable dimensions truly reflected specific visual or semantic properties, or whether they only superficially appeared to show this correspondence. To this end, we experimentally and causally manipulated images and observed the impact on dimension scores. Beyond general interpretability, these analyses further establish which visual features in each image drive individual dimensions and thus determine image representations.
@@
Image manipulation requires a direct mapping from input images to the embedding dimensions. Since the embedding dimensions were derived using a sampling based approach, this prohibits a direct link to manipulated or novel images. To overcome this challenge, we used $\ell_{2}$-regularized linear regression for linking penultimate layer activations of the DNN to each individual dimension of the learned embedding (Fig. 1d). Penultimate layer activations were indeed highly predictive of each embedding dimension, with all dimensions exceeding an $R^{2}$ of 75%, and the majority exceeding 85%. Thus, this allowed us to accurately predict dimension values for novel images.
@@
Having established an end-to-end mapping between input image and individual object dimensions, we next used three approaches to both probe the consistency of the interpretation and identify dimension-specific image features. First, to identify image regions relevant for each individual dimension, we used Grad-CAM (Selvaraju et al., 2017), an established technique for providing visual explanations. Grad-CAM generates heatmaps that highlight the image regions that are most influential for model predictions. Unlike traditional usage, which often focuses on creating visual explanations for model categorizations (e.g., dog vs. cat), we employed Grad-CAM to reveal what image regions drive the dimensions in the DNN embedding. The results of this analysis are illustrated with example images in Figure 3. Object dimensions were indeed driven by different image regions that contain relevant information, in line with the dimension’s interpretation derived from human ratings and suggesting that the representations captured by the DNN’s penultimate layer allow distinguishing between different object parts that carry different functional significance.
@@
Figure 4: Most exciting images for dimensions in our embedding. a, Using StyleGAN XL (Sauer et al., 2022), we optimized a latent code to maximize the predicted response in a specific embedding dimension. b, Visualizations for different dimensions in our embedding. We show the top $10$ images that score highest in the dimension and the corresponding top $10$ generated images.
@@
As a second image explanation approach, to specifically highlight what image features drive a dimension, we used a generative image model to create novel images optimized for maximizing values of a given dimension (Montavon et al., 2018; Yosinski et al., 2015; Erhan et al., 2009). Unlike conventional activation maximization that targets a single DNN unit or a cluster of units, our approach aimed to selectively amplify activation in dimensions of the DNN embedding across the entire DNN layer, using a pretrained generative adversarial neural network (StyleGAN XL; Sauer et al., 2022)). The results of this procedure are shown in Figure 4b. The approach successfully generated images with high scores in the dimensions of our DNN embedding. Indeed, the properties highlighted by these generated images appear to align with the common interpretation of each specific dimension, again suggesting that the DNN embedding contained conceptually meaningful and coherent object properties similar to those found in humans.
@@
Third, given that different visual features naturally co-occur across images, in order to tease apart their respective contribution, we causally manipulated individual image features and observed the effect on predicted DNN dimensions. We exemplify this approach with manipulations in color, object shape, and background (see Supplementary Information C). The results largely confirmed our predictions, leading to specific decreases or increases of activation in dimensions that appeared to be representing these features.
1.3 Factors underlying similarities and differences between humans and DNNs

### v2

@@ Abstract
Determining the similarities and differences between humans and artificial intelligence (AI) is an important goal both in computational cognitive neuroscience and machine learning, promising a deeper understanding of human cognition and safer, more reliable AI systems. Much previous work comparing representations in humans and AI has relied on global, scalar measures to quantify their alignment. However, without explicit hypotheses, these measures only inform us about the degree of alignment, not the factors that determine it. To address this challenge, we propose a generic framework to compare human and AI representations, based on identifying latent representational dimensions underlying the same behavior in both domains. Applying this framework to humans and a deep neural network (DNN) model of natural images revealed a low-dimensional DNN embedding of both visual and semantic dimensions. In contrast to humans, DNNs exhibited a clear dominance of visual over semantic properties, indicating divergent strategies for representing images. While in-silico experiments showed seemingly consistent interpretability of DNN dimensions, a direct comparison between human and DNN representations revealed substantial differences in how they process images. By making representations directly comparable, our results reveal important challenges for representational alignment and offer a means for improving their comparability.
Deep neural networks (DNNs) have achieved impressive performance, matching or surpassing human performance in various perceptual and cognitive benchmarks, including image classification (Krizhevsky et al., 2012; He et al., 2016), speech recognition (Hinton et al., 2012; Amodei et al., 2016) and strategic gameplay (Silver et al., 2016; Vinyals et al., 2019). In addition to their excellent performance as machine learning models, DNNs have drawn attention in the field of computational cognitive neuroscience for their notable parallels to cognitive and neural systems in humans and animal models (Khaligh-Razavi and Kriegeskorte, 2014; Yamins et al., 2014; Güçlü and van Gerven, 2015; Rajalingham et al., 2015; Kubilius et al., 2016). These similarities, observed through different types of behavior or patterns of brain activity, have sparked a growing interest in determining both the factors underlying these similarities and the differences between human and DNN representations. From the machine learning perspective, understanding where DNNs exhibit a limited alignment with humans can support the development of better and more robust artificial intelligence systems. From the perspective of computational cognitive neuroscience, DNNs with stronger human alignment promise to be better candidate computational models of human cognition and behavior (Cichy and Kaiser, 2019; Lindsay, 2021; Kanwisher et al., 2023; Doerig et al., 2023).
@@
Despite the appeal of summary statistics, such as correlation coefficients or explained variance, for comparing the representational alignment of DNNs with humans, they only quantify the degree of representational or behavioral alignment. However, without explicit hypotheses about potential causes for misalignment, such scalar measures are limited in their explanatory scope of which properties determine this degree of alignment, that is, which representational factors underlie the similarities and differences between humans and DNNs. While diverse methods for interpreting DNN activations have been developed at various levels of analysis, ranging from single units to entire layers (Erhan et al., 2009; Zeiler and Fergus, 2014; Zhou et al., 2018; Morcos et al., 2018; Bau et al., 2020), the direct comparability to human representations has remained a key challenge.
Inspired by recent work in the cognitive sciences that has revealed core visual and semantic representational dimensions underlying human similarity judgments of object images (Hebart et al., 2020), here we propose a framework to systematically analyze and compare the dimensions that shape representations in humans and DNNs. In this work, we apply this framework to human visual similarity judgments and representations in a DNN trained to classify natural images. Our approach reveals numerous interpretable DNN dimensions that appear to reflect both visual and semantic image properties and that appear to be well-aligned to humans. In contrast to humans, who showed a dominance of semantic over visual dimensions, DNNs exhibited a striking visual bias, demonstrating that downstream semantic behavior is driven more strongly by different, primarily visual, strategies. While psychophysical experiments on DNN dimensions underscored their global interpretability, a direct comparison with human dimensions revealed that DNN representations in fact only approximate human representations but lack the consistency expected from property-specific visual and semantic dimensions. Together, our results reveal key factors underlying the representational alignment and misalignment between humans and DNNs, shed light on potentially divergent representational strategies, and highlight the potential of this approach to identify the factors underlying the similarities and differences between humans and DNNs.
@@ 1 Results
To improve the comparability of human and DNN representations, we aimed to identify the similarities and differences in core dimensions underlying human and DNN representations of images. To achieve this aim, we treated the neural network analogously to a human participant carrying out a cognitive behavioral experiment and then derived representational embeddings using a recent variational embedding technique (Muttenthaler et al., 2022) both from human similarity judgments and a DNN on the same behavioral task (see Methods). This approach ensured direct comparability between human and DNN representations. As a behavioral task, we chose a triplet odd-one-out similarity task, where from a set of three object images $i$, $j$, $k$, participants have to select the most dissimilar object (Fig. 1a; see Supplementary Information D for an analysis of the role of task instructions on triplet choice behavior). In this task, the perceived similarity between two images $i$ and $j$ is defined as the probability of choosing these images to belong together across varying contexts imposed by a third object image $k$. By virtue of providing minimal contexts, the odd-one-out task highlights the information sufficient to capture the similarity between object images $i$ and $j$ across diverse contexts. In addition, it approximates human categorization behavior for arbitrary visual and semantic categories, even for quite diverse sets of objects (Zheng et al., 2019; Hebart et al., 2020; Muttenthaler et al., 2022). Thus, by focusing on the building blocks of categorization which underlies diverse behaviors, this task is ideally suited for comparing object representations between humans and DNNs.
Figure 1: Overview: A computational framework that captures core DNN object representations in analogy to humans by simulating behavioral decisions in an odd-one-out task. a, The triplet odd-one-out task, where a human participant or a DNN is presented a set of three images and is asked to select the image that is most different from the others. b, Sampling approach of odd-one-out decisions from DNN representations. First, a dot-product similarity space is constructed from DNN features. Next, for a given triplet of objects, the most similar pair in this similarity space is identified, making the remaining object the odd-one-out. For humans, this sampling approach is based on observed behavior, which is used as a measure of their internal cognitive representations. c, Illustration of the computational modeling approach to learn a lower-dimensional object representation for human participants and the DNN, optimized to predict behavioral choices made in the triplet task. d, Schematic depiction of the interpretability pipeline that allows for the prediction of object embeddings from pretrained DNN features.
@@
For human behavior, we used a set of 4.7 million publicly available odd-one-out judgments (Hebart et al., 2023) over 1,854 diverse object images, derived from the THINGS object concept and image database (Hebart et al., 2019). For the DNN, we collected similarity judgments for 24,102 images of the same objects used for humans (1,854 objects, 13 examples per object). We used a larger set of object images since the DNN was less limited by constraints in dataset size than humans. This allowed us to obtain more precise estimates of their representation. To derive DNN representations, we chose a pretrained VGG-16 model (Simonyan and Zisserman, 2014), given its common use in the computational cognitive neurosciences. Specifically, this network has been shown to exhibit good correspondence to both human behavior (Geirhos et al., 2018) and measured neural activity (Güçlü and van Gerven, 2015; Schrimpf et al., 2018; Nonaka et al., 2021) and performs well at predicting human similarity judgments (Jozwik et al., 2017; Peterson et al., 2018; King et al., 2019; Storrs et al., 2021; Muttenthaler et al., 2023a; Kaniuth et al., 2024). VGG-16 was trained on the 1000-class ImageNet dataset (Deng et al., 2009), which contains a diverse range of image categories, such as animals, everyday objects, and scenes. However, for completeness, we additionally ran similar analyses for a broader range of neural network architectures (see Supplementary Information A). We focused on the penultimate layer activations as they are closest to the behavioral output, and they also showed closest representational correspondence to humans (see Supplementary Information B). For the DNN, we generated a dataset of behavioral odd-one-out choices for the 24,102 object images (Fig. 1b). To this end, we first extracted the DNN layer activations for all images. Next, for a given triplet of activations ${\bm{z}}_{i}$, ${\bm{z}}_{j}$ and ${\bm{z}}_{k}$, we computed the dot product between each pair as a measure of similarity, then identified the most similar pair of images in this triplet, and designated the remaining third image as the odd-one-out. Given the excessively large number of possible triplets for all 24,102 images, we approximated the full set of object choices from a random subset of $20$ million triplets (Jain et al., 2016).
@@
From both sets of available triplet choices, we next generated two representational embeddings, one for humans and one for the DNN, where each embedding was optimized to predict the odd-one-out choices in humans and DNNs, respectively. In these embeddings, each object is described through a set of dimensions that represent interpretable object properties. To obtain these dimensions and for comparability to previous work in humans (Zheng et al., 2019; Hebart et al., 2020; Muttenthaler et al., 2022), we imposed sparsity and non-negativity constraints on the optimization, which support their interpretability and provide cognitively plausible criteria for dimensions(Hoyer, 2002; Murphy et al., 2012; Fyshe et al., 2015; Hebart et al., 2020, 2023). Sparsity constrained the embedding to consist of fewer dimensions, motivated by the observation that real-world objects are typically characterized by only a few properties. Non-negativity encouraged a parts-based description, where dimensions cannot cancel each other out. Thus, a dimension’s weight indicated its relevance in predicting an object’s similarity to other objects. During training, each randomly initialized embedding was optimized using a recent variational embedding technique (Muttenthaler et al., 2022) (for details, see Methods: Embedding optimization and pruning). The optimization resulted in two stable, low-dimensional embeddings, with 70 reproducible dimensions for the DNN embedding and 68 for the human embedding. The DNN embedding captured 84.03% of the total variance in image-to-image similarity, while the human embedding captured 82.85% of the total variance and 91.20% of the explainable variance given the empirical noise ceiling of the dataset.
1.1 DNN dimensions reflect diverse image properties
@@
Having identified stable, low-dimensional embeddings that are predictive of triplet odd-one-out judgments, we first assessed the interpretability of each identified DNN dimension by visualizing object images with large numeric weights. In addition to this qualitative assessment, we validated these observations for the DNN by asking 12 (6 female, 6 male) human participants to provide labels for each dimension separately (see Methods: Labeling dimensions and construction of word clouds). Similar to the core semantic and visual dimensions underlying odd-one-out judgments in humans described previously (Hebart et al., 2020; Muttenthaler et al., 2022; Hebart et al., 2023), the DNN embedding yielded many interpretable dimensions, which appeared to reflect both semantic and visual properties of objects. The semantic dimensions included taxonomic membership (e.g. food-related, technology-related, home-related) and other knowledge-related properties (e.g. softness), while the visual dimensions reflected visual-perceptual attributes (e.g. round, green, stringy), with some dimensions reflecting a composite of semantic and visual properties (e.g. green and organic) (Fig. 2a). Of note, the DNN dimensions also revealed a sensitivity to basic shapes, including roundness, boxiness and tube-shape. This suggests that, in line with earlier studies (Hermann et al., 2020; Singer et al., 2022), DNNs indeed learn to represent basic shape properties, an aspect that might not be apparent in their overt behavior (Geirhos et al., 2019).
@@
Despite the apparent similarities, there were, however, also striking differences found between humans and the DNN. First, overall, DNN dimensions were less interpretable than human dimensions, as confirmed by the evaluation of all dimensions by two independent raters (see Supplementary Information C). This indicates a global difference in how the DNN assigns images as being conceptually similar to each other. Second, while human dimensions were clearly dominated by semantic properties, many DNN dimensions were more visual-perceptual in nature or reflected a mixture of visual and semantic information. We quantified this observation by asking the same two independent experts to rate human and DNN dimensions according to whether they were primarily visual-perceptual, semantic, reflected a mixture of both, or were unclear (Fig. 2b). To confirm that the results were not an arbitrary byproduct of the chosen DNN architecture, we provided the raters with four additional DNNs for which we had computed additional representational embeddings. The results revealed a clear dominance of semantic dimensions in humans, with only a small number of mixed dimensions. In contrast, for DNNs we found a consistently larger proportion of dimensions that were dominated by visual information or that reflected a mixture of both visual and semantic information (Fig. 2c, Supplementary Fig. 1b for all DNNs). This visual bias is also present across intermediate representations of VGG-16 and even stronger in early to late convolutional layers (see Supplementary Fig. 2). This demonstrates a clear difference in the relative weight that humans and DNNs assign to visual and semantic information, respectively. We independently validated these findings using a semantic text embedding and observed a similar pattern of visual bias (see Supplementary Information E, indicating that the results were not solely a product of human rater bias).
@@ 1.2 Linking DNN dimensions to their interpretability
Figure 3: Relevance of image properties for embedding dimension. a, General methodology of the approach. We used Grad-CAM (Selvaraju et al., 2017) to visualize the importance of distinct image parts based on the gradients of the penultimate DNN features that we initially used to sample triplet choices. The gradients were obtained in our fully differentiable interpretability model with respect to a dimension ${\bm{w}}$ in our embedding. b, We visualize the heatmaps for three different images and dimensions. Each column shows the relevance of parts of an image for that dimension. For this figure, we filtered the embedding by images available in the public domain (Stoinski et al., 2023), except for two images sourced from Flickr under a CC BY 2.0 license: the flashlight by Borysiuk (2010) and the wineglass by Szkutnik (2012).
@@
Despite the overall differences in human and DNN representational dimensions, the DNN also contained many dimensions that appeared to be interpretable and comparable to those found in humans. Next, we aimed at testing to what degree these interpretable dimensions truly reflected specific visual or semantic properties, or whether they only superficially appeared to show this correspondence. To this end, we experimentally and causally manipulated images and observed the impact on dimension scores. Beyond general interpretability, these analyses further establish which visual properties in each image drive individual dimensions and thus determine image representations.
@@
Image manipulation requires a direct mapping from input images to the embedding dimensions. However, the embedding dimensions were derived using a sampling-based optimization based on odd-one-out choices inferred from penultimate DNN features. Consequently, this approach does not directly map these features to the learned embedding. To establish this mapping, we applied $\ell_{2}$-regularized linear regression to link the DNN’s penultimate layer activations to the learned embedding. This mapping then enables the prediction of embedding dimensions from the penultimate features activations in response to novel or manipulated images. (Fig. 1d). Penultimate layer activations were indeed highly predictive of each embedding dimension, with all dimensions exceeding an $R^{2}$ of 75%, and the majority exceeding 85%. Thus, this allowed us to accurately predict dimension values for novel images.
@@
Having established an end-to-end mapping between input image and individual object dimensions, we next used three approaches to both probe the consistency of the interpretation and identify dimension-specific image properties. First, to identify image regions relevant for each individual dimension, we used Grad-CAM (Selvaraju et al., 2017), an established technique for providing visual explanations. Grad-CAM generates heatmaps that highlight the image regions that are most influential for model predictions. Unlike the typical use of Grad-CAM, which focuses on generating visual explanations for model classifications (e.g., dog vs. cat), we employed Grad-CAM to reveal which image regions drive the dimensions in the DNN embedding. The results of this analysis are illustrated with example images in Fig. 3. Object dimensions were indeed driven by different image regions that contain relevant information, in line with the dimension’s interpretation derived from human ratings and suggesting that the representations captured by the DNN’s penultimate layer allow us to distinguish between different parts of the image that carry different functional significance.
@@
Figure 4: Maximally activating images for embedding dimensions. a, Using StyleGAN XL (Sauer et al., 2022), we optimized a latent code to maximize the predicted response in a specific embedding dimension. b, Visualizations for different dimensions in our embedding. We show the top $10$ images that score highest in the dimension and the corresponding top $10$ generated images. For this figure, we filtered the embedding by images available in the public domain (Stoinski et al., 2023)
@@
As a second image explanation approach, to highlight which image properties drive a dimension, we used a generative image model to create novel images optimized for maximizing values of a given dimension (Erhan et al., 2009; Yosinski et al., 2015; Montavon et al., 2018). Unlike conventional activation maximization targeting a single DNN unit or a cluster of units, our approach aimed to selectively amplify activation in dimensions of the DNN embedding across the entire DNN layer, using a pretrained generative adversarial neural network (StyleGAN XL; Sauer et al., 2022)). To achieve this, we applied our linear end-to-end mapping to predict the embedding dimensions from the penultimate activations in response to the images generated by StyleGAN-XL. The results of this procedure are shown in Figure 4b. The approach successfully generated images with high numerical values in the dimensions of our DNN embedding. Indeed, the properties highlighted by these generated images appear to align with human assigned labels for each specific dimension, again suggesting that the DNN embedding contained conceptually meaningful and coherent object properties similar to those found in humans.
@@
As a third image explanation approach, given that different visual properties naturally co-occur across images, and to tease apart their respective contribution, we causally manipulated individual image properties and observed the effect on predicted DNN dimensions. We exemplify this approach with manipulations in color, object shape, and background (see Supplementary Information F), largely confirming our predictions, showing specific activation decreases or increases in dimensions that appeared to be representing these properties.
1.3 Factors driving human and DNN similarities and differences